Instructions: Questions labeled with (R) require use of R. Please provide appropriate R commands and their output, along with sufficient explanation and interpretation of the output to demonstrate your understanding. All other questions should be completed without reference to any R commands or output. Make sure you give enough explanation so your tutor can follow your reasoning if you happen to make a mistake. Please also try to be as succinct as possible. Each assignment will include marks tor good presentation and for attempting all problems1. (R) An interior automotive supplier places several electrical wires in a harness. A pull test measures the force required to pull spliced wires apart. A customer requires that each wire that is spliced into the harness withstand a pull force of 20 pounds. Let X be a random variable representing the pull force required to pull a spliced wire apart. The following are 20 observations of X 28.8 24.4 30.1 25.6 26.4 23.9 22.1 22.5 27.6 28.1 20.8 27.7 24.4 25.1 24.6 26.3 28.2 22.2 26.3 24.4 (a) Give basic summary statistics for these data and produce a box plot. Briefly comment on center, spread and shape of the distribution. (b) Assuming a normal distribution, compute maximum likelihood estimates for the parameters. (c) Draw a density histogram and superimpose a pdf for a normal distribution using the estimated parameters. (d) Draw a QQ plot to compare the data against the fitted normal distribution Include a reference line. Comment on the fit of the model to the data

NOTE: PLEASE INCLUDE ALL THE R CODES ALONG WITH THE PLOTS
Answer #1

(a)

R code:

x=c(28.8,24.4,30.1,25.6,26.4,23.9,22.1,22.5,27.6,28.1,20.8,27.7,24.4,25.1,24.6,
26.3,28.2,22.2,26.3,24.4)
summary(x)
boxplot(x)

Output:

Min. 1st Qu. Median Mean 3rd Qu. Max.
20.80 24.28 25.35 25.48 27.63 30.10

26 30 24 26 2

Center=25.35, Inter quartile range=27.63-24.28=3.35. Since 3rd quartile-2nd quartile>2nd quartile-1st quartile so the distribution is positively skewed.

(b)~Suppose~X_1,X_2,...,X_n~are~random~sample~from~N(mu,sigma^2). The~likelihood~function=L(mu,sigma^2)=prod_{i=1}^nrac{1}{sqrt{2pi}sigma}expleft(-rac{1}{2sigma^2}(x_i-mu)^2 ight ) =rac{1}{(2pisigma^2)^{rac{n}{2}}}expleft(-rac{1}{2sigma^2}sum_{i=1}^n(x_i-mu)^2 ight ) l(mu,sigma^2)=log L(mu,sigma^2)=-rac{n}{2}log 2pi-rac{n}{2}log sigma^2-rac{1}{2sigma^2}sum_{i=1}^n(x_i-mu)^2 rac{partial l(mu,sigma^2)}{partial mu}=rac{1}{sigma^2}sum_{i=1}^n(x_i-mu)=0Rightarrow hat{mu}=rac{1}{n}sum_{i=1}^nx_i=ar{x} rac{partial l(mu,sigma^2)}{partial sigma^2}=-rac{n}{2sigma^2}+rac{1}{2sigma^4}sum_{i=1}^n(x_i-mu)^2=0Rightarrow hat{sigma}^2=rac{1}{n}sum_{i=1}^n(x_i-hat{mu})^2 =rac{1}{n}sum_{i=1}^n(x_i-ar{x})^2

0o is negative definite then

Hence M LEs of μ and σ2 are i and-y (ri-i)2. 7t For this data set, ML estimates are: r = 25 475 ,-〉 (xi-x)2-5.9069

(c)

x=c(28.8,24.4,30.1,25.6,26.4,23.9,22.1,22.5,27.6,28.1,20.8,27.7,24.4,25.1,24.6,
26.3,28.2,22.2,26.3,24.4)
h<-hist(x, breaks=6, col="red", xlab="Pull force",
main="Histogram with Normal Curve",ylim=c(0,7))
xfit<-seq(min(x),max(x),length=40)
yfit<-dnorm(xfit,mean=mean(x),sd=sqrt(5.9069))
yfit <- yfit*diff(h$mids[1:2])*length(x)
lines(xfit, yfit, col="blue", lwd=2)

Histogram with Normal Curve LO しの CN 20 24 26 28 30 32 Pu force

(d)

R code:

x=c(28.8,24.4,30.1,25.6,26.4,23.9,22.1,22.5,27.6,28.1,20.8,27.7,24.4,25.1,24.6,
26.3,28.2,22.2,26.3,24.4)
qqnorm(x,lwd=2);qqline(x, col = 2,lwd=2)

Normal Q-Q Plot 0 CH 丈 CH 2 0 1 Theoretical Quantiles

It is observed that the sample quantiles are closed to reference line hence we can assume data come from normal distribution.